John is 4 years younger than her brother. Five years from now, the sum of their ages will be 32. Find John's present age.
Answer by Alan3354(21599)

John is 4 years younger than her brother. Five years from now, the sum of their ages will be 32. Find John's present age.
-------------------
John's a girl?
J = John's age
B = brother's age
----------------
John is 4 years younger than her brother.
So, B = J + 4 eqn 1
--------------
Five years from now, the sum of their ages will be 32.
(J+5) + (B+5) = 32
J + B = 22 eqn 2
Substitute for B from eqn 1
J + (J + 4) = 22
2J + 4 = 22
2J = 18
J = 9
B = 9+4
B = 13